For a certain sporting item that is produced in a factory, we have that the income is given by the function I (x) = x (800-4x), while the cost of production is given by C (x) = 0.4x2 + 8x + 800, where x are the units produced.

The marginal income from the sale and production of article 50 is:

 [Value 2 points]

Select one:
a. 600
b. 3000
c. 2200
d. 400
Answer #1

Answer :

The income function is given by :

I (x) = x * (800 - 4x) ; where x = number of units produced

Or, I (x) = 800x - 4x^2

Marginal income (x) = Differentiating the income function with respect to x we get = 800 - (2 * 4x) = 800 - 8x

When x = 50

Marginal income (x = 50) = 800 - 8*50 = 800 - 400 = 400

Answer : Option d) 400